OpenStudy (anonymous):

Simplify the following expression completely. (5 x -10)/(45) Enter the numerator and denominator separately in the boxes below. If the denominator is 1, enter the number 1. Do not leave either box blank. Answer:

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

5x - 10 _______ 45 is that the equation? Sorry I couldn't read it right

OpenStudy (anonymous):

yes that's it @jadeybells

OpenStudy (anonymous):

\[\huge \frac{(5 x -10)}{(45)} \] From numerator we can see that 5 is common in both the terms, hence we take it common. \[\huge = \frac{5(x -2)}{ 9 \times 5} \] now cancel out 5 from Nr. and Dr. we find \[\huge = \frac{(x -2)}{ 9} \] @cancherolugano